OpenStudy (anonymous):

Solve for x: (x - 9)2 = 1

OpenStudy (anonymous):

take the square root of both sides x - 9 = 1 or x - 9 = -1 Add 9 to both sides x = 10 or x = 8

OpenStudy (anonymous):

X= 1by2 + 9

OpenStudy (anonymous):

Divide both sides by 2 and you get x-9= 1/2. Now add 9 to both sides, giving you x= 1/2 +9. 9 plus 1/2 = 9.5 Therefore x=9.5